Thatobaked 25 muffins. He gave 3/25 of the muffins to his friend and 9/25 of the muffins to his sister.what fraction of the muffins is left?


1
Expert's answer
2022-02-11T09:28:51-0500

Explanation & Calculations


  • The math involved here has been made easy by making the denominator of each fraction equal to the total number of muffins.
  • So you can simply think of the distribution as 3 out of 25 and 9 out of 25 and after all, (3+9 = ) 12 out of 25 muffins would have been distributed away.
  • So, (25-12 = )13 muffins will be remaining after it.
  • Therefore, the fraction of the remaining muffins is "\frac{13}{25}"